OpenStudy (anonymous):

what is this equation 1/2*m*v^2 = 3/2*k*T?

OpenStudy (anonymous):

Kinetic Energy =???

OpenStudy (anonymous):

Kinetic Energy

OpenStudy (anonymous):

what is the purpose of the equation? what does the equation describe? Ive never seen 3/2*k*t

OpenStudy (stormfire1):

It's used to calculate the kinetic energy of an ideal gas where k is Boltzmann's constant and T is the temperature of the gas in Kelvin. For a better explanation of the equation go to http://www.ux1.eiu.edu/~cfadd/1150/14Thermo/kinetic.html

OpenStudy (stormfire1):

Absolutely nothing would happen...literally :)

OpenStudy (stormfire1):

Note the intended pun.

OpenStudy (anonymous):

seriously?? what would happen based on the equation?

OpenStudy (stormfire1):

\[\frac{1}{2}mv^2=\frac{3}{2}kT\]so if T was 0:\[\frac{1}{2}mv^2=\frac{3}{2}k*0\]which to me implies what we already know...velocity must also be 0.

OpenStudy (stormfire1):

If you're asking what would happen if you went below absolute 0 then I'd say the question is flawed. Absolute 0 is the idea that all particles essentially have 0 velocity...

OpenStudy (anonymous):

@shakir

OpenStudy (anonymous):

@ajprincess

OpenStudy (anonymous):

@Brent0423 it is kinetic energy equation and k is called boltzman constant,

OpenStudy (anonymous):

KE=3/2KT

OpenStudy (anonymous):

@shakir what do u think?

OpenStudy (anonymous):

@stormfire1 is correct @Brent0423 .....

OpenStudy (ajprincess):

I too agree with @stomfire1's answer. @Brent0423

OpenStudy (anonymous):

@shakir